OpenStudy (anonymous):

Simplify the rational expression. State any excluded values. x^2-36/6-x

OpenStudy (anonymous):

a^2 - b^2 = (a-b)(a+b) can you do something with the numerator now ?

OpenStudy (anonymous):

im just confused im not really good at math

OpenStudy (anonymous):

ok so using : a^2 - b^2 = (a-b)(a+b) we can say a^2 = x^2 and b^2 = 36 take only positive values so it means that b = 6 and a = x so.. what is (a-b)(a+b) ?